OpenStudy (bohotness):

Which ordered pairs are solutions to the equation? y = 2x + 5 A. (5, 15) and (9, 2) B. (15, 5) and (2, 9) C. (5, 10) and (2, 9) D. (5, 15) and (2, 9)

OpenStudy (dtan5457):

They have 2 coordinate points. Plug the x value of each one and see if the corresponding y value comes out correctly. For example (5,15) plugged into your function of y=2x+5 15=2(5)+5 SO (5,15) works..

OpenStudy (dtan5457):

Just returned after 2 weeks of not using the site, only to see another direct answer. Sigh, I should just quit.

OpenStudy (bohotness):

I am here oh

OpenStudy (bibby):

DO IT

OpenStudy (bohotness):

@ilovewolf

OpenStudy (bibby):

lazy lazy

OpenStudy (ilovewolf):

bibby stop being mean

OpenStudy (nincompoop):

That is one way to look for the "right answer" but there's another way that involves actually solving it One must remember that in the slope-intercept form, \(y = mx+b \), the \(b \) is the y-intercept. That is when the line touches and passes through the y-axis. Graphically, it would look like this:

OpenStudy (nincompoop):

Assuming that once the y-intercept is identified, \((x =0, y = b) \), we can further solve for another coordinate easily, that is to know our x-intercept, that is when the value of \(y=0 \) the line touches and passes through the x-axis. Graphically it is represented by this: 

OpenStudy (nincompoop):

So if we already know that the value of y is zero, we can solve analytically by setting our equation to zero and solve for x, which will give us something like this: \(0= mx + b \)

OpenStudy (nincompoop):

It is one thing to figure out the answer by trial and error, but that doesn't show if you understood anything about linear equations at all. Through display of graphs and analysis, it would show the amount of information you actually know and able to use.